OpenStudy (anonymous):

Which literal equation is equivalent to d = 0.2t?

OpenStudy (anonymous):

A.t=d/0.2 B.t=0.2/d

OpenStudy (anonymous):

C.\[0.2=\frac{ t}{ d }\] D.\[0.2=td\]

OpenStudy (anonymous):

@Compassionate @damara1213 @Preetha @danish071996 @mathmath333

OpenStudy (anonymous):

d = 0.2t FIRST SCENARIO, SOLVE FOR t SECOND SCENARIO, SOLVE FOR 0.2 NOW COMPARE YOUR ANSWERS TO THE GIVEN CHOICES

OpenStudy (anonymous):

?how do i solve for them

OpenStudy (anonymous):

that what i got i dont get it eather sorry

OpenStudy (anonymous):

its a